### v1

# Spectral Concentration at the Edge of Stability: Information Geometry of Kernel Associative Memory

[…]
Spectral Concentration at the Edge of Stability: Information Geometry of Kernel Associative Memory
keywords:
Kernel Hopfield Network, Information Geometry, Edge of Stability, Spectral Concentration, Self-Organized Criticality
{summary}
High-capacity kernel Hopfield networks exhibit a “Ridge of Optimization” characterized by extreme stability. While previously linked to ”Spectral Concentration,” its origin remains elusive. Here, we analyze the network dynamics on a statistical manifold, revealing that the Ridge corresponds to the “Edge of Stability,” a critical boundary where the Fisher Information Matrix becomes singular. We demonstrate that the apparent Euclidean force antagonism is a manifestation of Dual Equilibrium in the Riemannian space. This unifies learning dynamics and capacity via the Minimum Description Length principle, offering a geometric theory of self-organized criticality.
1 Introduction
Associative memory is fundamentally a geometric problem: how to embed discrete patterns as stable fixed points within a continuous state space. Recent advances using Kernel Logistic Regression (KLR) have demonstrated that learning can sculpt these landscapes to achieve capacities far exceeding classical limits [1, 2, 3]. Our previous phenomenological analysis identified a “Ridge of Optimization” where stability is maximized via a mechanism we termed Spectral Concentration—a state where the weight spectrum exhibits a sharp hierarchy [4].
However, a deeper question remains: Why does the learning dynamics self-organize into this specific spectral state? Why does the system operate at the brink of instability? To answer these questions, we must look beyond the Euclidean geometry of the weight parameters and consider the intrinsic geometry of the probability distributions they represent. This is the domain of Information Geometry [5].
In this work, we reinterpret the KLR Hopfield network as a statistical manifold equipped with a Fisher-Rao metric. We show that:
  1. 1.
    The “Ridge” is the locus where the Fisher Information Matrix (FIM) undergoes a spectral phase transition, effectively minimizing the information dimension of the memory.
  2. 2.
    The optimization dynamics on the Ridge satisfies a Dual Balance, where the massive Euclidean gradients are neutralized by the curvature of the manifold, resulting in a stationary state in the dual affine coordinates.
Although our focus is on associative memory, the geometric mechanisms revealed here, specifically Spectral Concentration and Dual Equilibrium, may offer broader insights into the generalization capabilities of over-parameterized neural networks, where similar phenomena of “feature learning” and “sharpness selection” have been reported.
2 Geometric Framework
We consider the kernel Hopfield network not merely as a dynamical system, but as a parametric statistical model.
2.1 Statistical Manifold of Kernel Memory
Let $\mathcal{H}$ be a reproducing kernel Hilbert space with kernel $K(\cdot,\cdot)$. The state of the network is determined by the dual variables $\bm{\alpha}\in\mathbb{R}^{P\times N}$, where $P$ is the number of stored patterns and $N$ is the number of neurons. For a single neuron $i$, the probability of being in state $s_{i}=+1$ given the input pattern $\bm{\xi}^{\mu}$ is modeled by the logistic sigmoid function:
$p(\bm{\xi}^{\mu};\bm{\alpha}_{i})=\sigma\left(\sum_{\nu=1}^{P}\alpha_{\nu i}K(\bm{\xi}^{\mu},\bm{\xi}^{\nu})\right).$ (1)
The set of all such realizable probability distributions forms a statistical manifold $\mathcal{M}=\{p(\cdot;\bm{\alpha})\mid\bm{\alpha}\in\mathbb{R}^{P\times N}\}$.
2.2 Fisher Information Matrix (FIM)
The intrinsic geometry of $\mathcal{M}$ is governed by the Fisher Information Matrix (FIM) $G(\bm{\alpha})$. For the weights associated with a single neuron (omitting index $i$), the FIM is given by:
$G_{\mu\nu}(\bm{\alpha})=\mathbb{E}\left[\frac{\partial\log p}{\partial\alpha_{\mu}}\frac{\partial\log p}{\partial\alpha_{\nu}}\right].$ (2)
In the context of KLR, this takes the specific form involving the kernel Gram matrix $\bm{K}$:
$G(\bm{\alpha})=\bm{K}\bm{D}(\bm{\alpha})\bm{K},$ (3)
where $\bm{D}(\bm{\alpha})$ is a diagonal matrix with entries $D_{\mu\mu}=p_{\mu}(1-p_{\mu})$ representing the variance of the prediction for pattern $\mu$.
2.3 Riemannian vs. Euclidean Gradients
The learning dynamics can be viewed as a flow on this manifold. The standard gradient descent follows the steepest direction in the Euclidean metric, $\Delta\bm{\alpha}\propto-\nabla_{\bm{\alpha}}L$. However, the intrinsic steepest direction is given by the Natural Gradient [6], which accounts for the curvature $G$:
$\tilde{\nabla}L=G^{-1}\nabla_{\bm{\alpha}}L.$ (4)
Our central hypothesis is that the “Ridge” phenomena emerge from the interplay between these two vector fields—the Euclidean force $\nabla L$ and the geometric curvature $G$.
3 Spectral Concentration as Information Compression
Our previous phenomenological study identified “Spectral Concentration” in the weight matrix $\bm{\alpha}$ as the key to stability. Here, we show that this is a direct consequence of the information geometry of the learned manifold.
3.1 Effective Dimensionality of Memory
The rank of the Fisher Information Matrix $G$ determines the number of independent directions in the probability space that the model can represent locally. We analyze the eigenvalue spectrum of $G$, denoted by $\lambda_{1}\geq\lambda_{2}\geq\dots\geq\lambda_{P}$. We define the Effective Information Dimension $d_{\text{eff}}$ using the stable rank:
$d_{\text{eff}}(G)=\left(\sum_{k=1}^{P}\lambda_{k}\right)^{2}\left(\sum_{k=1}^{P}\lambda_{k}^{2}\right)^{-1}.$ (5)
Observation 1 (L-Shaped Spectrum): Numerical analysis reveals Fig. 1 reveals a striking contrast in the spectral structures. In the local regime (blue dashed line), the spectrum is flat ($\lambda_{k}\approx\lambda_{1}$), indicating a diffuse distribution of information where no dominant direction exists. In contrast, on the Ridge (red solid line), the spectrum exhibits a characteristic “L-shape”: the normalized eigenvalues drop by orders of magnitude immediately after the leading mode ($\lambda_{2}\ll\lambda_{1}$), yet the tail remains non-zero and flat ($\lambda_{k>1}>0$). This confirms that the Ridge achieves stability via Spectral Concentration—compressing noise into a single dominant mode while preserving the dimensionality required for memory capacity. Quantitatively, while the local regime exhibits $d_{\text{eff}}(G)\approx P$ (full utilization of dimensions), the Ridge collapses this to $d_{\text{eff}}(G)\approx 1.5\sim 5.0$, depending on the load. This drastic reduction confirms that the network performs optimal compression, retaining only the essential degrees of freedom.
Figure 1: Spectrum of the Fisher Information Matrix $G$ (normalized by $\lambda_{1}$). Blue (Off Ridge, Local): The spectrum is flat ($\lambda_{k}/\lambda_{1}\approx 1$), indicating a lack of structural compression. Red (On Ridge): The spectrum exhibits Spectral Concentration, dropping sharply after the first mode but maintaining a non-zero tail. This “L-shaped” profile signifies optimal information compression. Green (Off Ridge, Global): The tail collapses to near zero, indicating rank deficiency.
3.2 Minimum Description Length (MDL) Principle
This spectral structure can be interpreted through the Minimum Description Length (MDL) principle [7]. The network compresses the high-dimensional input noise into a single dominant mode ($\lambda_{1}$) to maximize robustness, while maintaining just enough degrees of freedom ($\lambda_{k>1}$) to separate the $P$ stored patterns. Unlike a Rank-1 collapse ($G\to\bm{v}_{1}\bm{v}_{1}^{\top}$), where pattern distinguishability is lost ($\lambda_{k>1}\to 0$), the Ridge maintains a “working margin” of information dimensionality.
4 The Edge of Stability
Why does the learning dynamics stop at this state of Spectral Concentration? We propose that the Ridge corresponds to the limit of stability for the optimization dynamics.
4.1 Maximal Curvature at the Ridge
The local curvature of the statistical manifold is quantified by the largest eigenvalue of the FIM, $\lambda_{\max}(G)$. In gradient descent dynamics, the stability of the update steps is constrained by this curvature; excessively sharp curvature can destabilize the learning process (Edge of Stability [8]).
Figure 2: Phase diagram of the maximum eigenvalue of the Fisher Information Matrix, $\log_{10}(\lambda_{\max}(G))$. The Ridge (bright diagonal band) corresponds to the region where the manifold curvature is locally maximized ($\lambda_{\max}\gg 1$), effectively creating the “Edge of Stability.” In contrast, the dark region (bottom right) indicates information collapse due to saturation of the sigmoid function ($p\to 0/1$), leading to vanishing curvature.
Observation 2 (Criticality): Our phase diagram analysis (Fig. 2) reveals that $\lambda_{\max}(G)$ is maximized exactly on the Ridge. While in the local regime (top region), the curvature is relatively small ($\log_{10}\lambda_{\max}\approx-1.2$), on the Ridge, it reaches values orders of magnitude higher ($\log_{10}\lambda_{\max}\approx 1.5$), indicating the formation of extremely sharp attractors. Crucially, immediately beyond the Ridge (in the global regime, bottom right), the value drops precipitously ($\log_{10}\lambda_{\max}<-8.0$). This corresponds to a phase where the learning saturates ($p_{\mu}\to 0$ or $1$), causing the Fisher information to vanish due to the vanishing gradients of the sigmoid function. Thus, the Ridge represents the critical boundary where the system maximizes geometric curvature to the limit allowed before information collapse occurs.
4.2 The Cliff of Optimization
Thus, the “Ridge of Optimization” can be geometrically defined as the Edge of Stability: the manifold boundary where the curvature is maximal.
$\text{Ridge}\approx\left\{(\gamma,P/N)\mid\lambda_{\max}(G)\text{ is locally maximized}\right\}.$
The network self-organizes to this cliff edge because sharper attractors (higher curvature) provide stronger error correction. The system climbs the curvature gradient until it reaches the limit imposed by the learning dynamics or the saturation of the nonlinearity.
5 Dual Geometry and Force Antagonism
Our previous work identified a phenomenological “Force Antagonism” on the Ridge: the Direct Force $\bm{F}_{d}$ and Indirect Force $\bm{F}_{i}$ become massive and strongly anti-correlated ($\bm{F}_{d}\approx-\bm{F}_{i}$). We now reveal the geometric origin of this antagonism.
5.1 Euclidean vs. Riemannian Forces
The learning dynamics can be analyzed in two distinct metric spaces:
  1. 1.
    Euclidean Space ($\mathbb{R}^{P}$): The standard parameter space where weights $\bm{\alpha}$ reside. The driving force is the Euclidean gradient $\nabla L$.
  2. 2.
    Riemannian Space ($\mathcal{M}$): The statistical manifold equipped with the Fisher metric $G$. The intrinsic driving force is the Natural Gradient $\tilde{\nabla}L=G^{-1}\nabla L$.
Observation 3 (Dual Balance): Numerical experiments (Fig. 3) demonstrate a striking contrast on the Ridge:
  • The Euclidean gradient norm $\|\nabla L\|^{2}$ diverges (bright region), indicating massive internal stresses.
  • The Riemannian gradient norm $\|\nabla L\|_{G^{-1}}^{2}=\nabla L^{\top}G^{-1}\nabla L$ is minimized (dark region), indicating that the system is close to an intrinsic equilibrium.
(a): Euclidean Gradient Norm
(b): Riemannian Gradient Norm
Figure 3: Comparison of Gradient Norms: (a) Euclidean Gradient $\|\nabla L\|^{2}$ vs. (b) Riemannian (Natural) Gradient $\|\tilde{\nabla}L\|_{G^{-1}}^{2}$. On the Ridge, the Euclidean norm diverges (bright band in a), reflecting the massive “Force Antagonism.” However, the Riemannian norm is minimized (dark band in b), revealing that the system is actually in a state of Dual Equilibrium with respect to the intrinsic geometry. This confirms that the apparent instability in parameter space is resolved by the curvature of the statistical manifold.
5.2 Geometric Derivation of Antagonism
We define the “Indirect Force” in the language of information geometry as the correction term arising from the curvature of the manifold. The Natural Gradient relation $\nabla L=G\tilde{\nabla}L$ implies that the observed Euclidean force is the intrinsic force distorted by the metric $G$. On the Ridge, we found that the spectrum of $G$ is dominated by a leading eigenvalue $\lambda_{1}\gg 1$ with eigenvector $\bm{v}_{1}$. Approximating $G\approx\lambda_{1}\bm{v}_{1}\bm{v}_{1}^{\top}$, the relationship becomes:
$\nabla L\approx\lambda_{1}(\bm{v}_{1}^{\top}\tilde{\nabla}L)\bm{v}_{1}.$ (6)
This explains the “amplification” mechanism: even a small intrinsic mismatch ($\tilde{\nabla}L$) is magnified by the huge curvature $\lambda_{1}$, resulting in a massive Euclidean gradient $\nabla L$.
The decomposition $\nabla L=\bm{F}_{d}+\bm{F}_{i}$ can be understood as splitting the gradient into the “signal” component and the “curvature correction” component. The condition $\nabla L\to\bm{0}$ (equilibrium) in Euclidean space implies $\bm{F}_{d}\approx-\bm{F}_{i}$. However, the condition $\tilde{\nabla}L\to\bm{0}$ in Riemannian space is much stronger and more fundamental. The Ridge represents a state where the system has found a Dual Equilibrium: a point where the intrinsic error is minimized despite the extreme curvature of the ambient space.
Interestingly, the massive curvature $\lambda_{1}$ on the Ridge implies that the Natural Gradient step in the principal direction, scaled by $\lambda_{1}^{-1}$, becomes negligible. This suggests a self-stabilizing mechanism: as the system approaches the Edge of Stability, the intrinsic update step vanishes, naturally braking the learning dynamics exactly at the optimal boundary.
6 Discussion
In this study, we have reinterpreted the dynamics of KLR-trained Hopfield networks through the lens of Information Geometry. Our analysis reveals that the phenomenological “Ridge of Optimization” is not merely a sweet spot of hyperparameters, but a manifestation of fundamental geometric principles: Spectral Concentration, Edge of Stability, and Dual Equilibrium.
6.1 Self-Organized Criticality in Memory
Our finding that the optimal memory state lies at the brink of information collapse (Fig. 2) aligns with the hypothesis of Self-Organized Criticality (SOC) [9] in biological neural networks. The brain is thought to operate at a critical point to maximize information transmission and dynamic range. Similarly, our KLR Hopfield network tunes itself to the “Edge of Stability,” maximizing the curvature of the statistical manifold to create deep attractors, while stopping just short of the singular region where memory capacity vanishes. This suggests that SOC may be a universal requirement for high-capacity associative memory.
6.2 Implications for Deep Learning Generalization
The phenomenon of Spectral Concentration may offer broader insights into the generalization mechanics of over-parameterized neural networks. Recent studies in deep learning have reported a “feature learning” regime, often referred to as Neural Collapse [10], where the weight spectrum becomes low-rank to capture task-relevant structure. Our results provide a geometric explanation for this: the network minimizes the effective dimensionality (via the Fisher Information spectrum) to ensure robust inference (Dual Equilibrium), effectively performing a geometric Occam’s razor.
6.3 The Geometry of Attention Mechanisms
Modern Hopfield Networks are mathematically equivalent to the attention mechanism in Transformers [11]. Our geometric framework implies that the effectiveness of attention layers may also rely on specific spectral structures in their Key-Query matrices. If attention heads also operate at an “Edge of Stability” to sharpen their focus, our duality theory (Euclidean force vs. Riemannian curvature) could provide a new tool for analyzing and stabilizing the training of large-scale Transformer models. Future work will investigate whether trained Transformers exhibit similar spectral signatures to those found on the Ridge.
7 Conclusion
This geometric framework unifies the apparently contradictory observations of instability (high curvature, Edge of Stability) and stability (deep attractors, Dual Balance). The “Ridge of Optimization” is not merely a parameter sweet spot, but a manifestation of Self-Organized Criticality in the information geometry of neural networks. The network tunes itself to the brink of instability, maximizing curvature to carve the deepest possible attractor basins, while maintaining dual equilibrium to ensure learning convergence.